# Second Order Bounds for Contextual Bandits with Function Approximation

### v1

@@ Abstract
Many works have developed algorithms no-regret algorithms for contextual bandits with function approximation, where the mean rewards over context-action pairs belongs to a function class $\mathcal{F}$. Although there are many approaches to this problem, one that has gained in importance is the use of algorithms based on the optimism principle such as optimistic least squares. It can be shown the regret of this algorithm scales as $\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathrm{eluder}}(\mathcal{F})\log(\mathcal{F})T}\right)$ where $d_{\mathrm{eluder}}(\mathcal{F})$ is a statistical measure of the complexity of the function class $\mathcal{F}$ known as eluder dimension. Unfortunately, even if the variance of the measurement noise of the rewards at time $t$ equals $\sigma_{t}^{2}$ and these are close to zero, the optimistic least squares algorithm’s regret scales with $\sqrt{T}$. In this work we are the first to develop algorithms that satisfy regret bounds of the form $\widetilde{\mathcal{O}}\left(d_{\mathrm{eluder}}(\mathcal{F})\sqrt{\log(\mathcal{F})\sum_{t=1}^{T}\sigma_{t}^{2}}+d_{\mathrm{eluder}}(\mathcal{F})\cdot\log(|\mathcal{F}|)\right)$ for contextual bandits with function approximation and unknown variances. These bounds generalize existing techniques for deriving second order bounds in contextual linear problems.
@@ 1 Introduction
The focus of many works, including this one is to flesh out the consequences of different modeling assumptions governing the relationship between the context, the action and the reward; one of the most salient features of this model. For example by developing algorithms for scenarios where the reward is a linear function of a linear function of an embedding of the context and action pair [6, 33, 11, 1]. This has lead to algorithms such as OFUL that can be used to derive bounds for contextual bandit problems with linear rewards [1]. Other works have considered scenarios that go beyond the linear case, where it is assumed the reward function over context action pair $x_{t},a_{t}$ is realized by an unknown function $f_{\star}$ belonging to a known function class $\mathcal{F}$ (which can be more complex than linear). Various adaptive learning procedures compatible with generic function approximation have been proposed for contextual bandit problems. Among these, we highlight two significant methods relevant to our discussion; the Optimistic Least Squares algorithm introduced by [34] and the SquareCB algorithm introduced by [14]. Both of these methods offer guarantees for cumulative regret. Specifically, the cumulative regret of Optimistic Least Squares scales with factors $\mathcal{O}(\sqrt{d_{\mathrm{eluder}}(\mathcal{F})\log(|\mathcal{F}|)})$, while the cumulative regret of Square CB scales as $\mathcal{O}\left(\sqrt{|\mathcal{A}|\log(|\mathcal{F}|)}\right)$, where $\mathcal{A}$ is the set of actions. The eluder dimension [^1] We formally introduce this quantity in Section 2. Here we use a simpler notation to avoid confusion. ($d_{\mathrm{eluder}}$) is a statistical complexity measure introduced by [34], that enables deriving guarantees for adaptive learning algorithms based on the principle of optimism in contextual bandits and reinforcement learning [25, 18, 31, 10].
@@ 3 Optimistic Least Squares
The algorithms we propose in this work are based on the optimism principle. This simple yet powerful algorithmic idea is the basis of a celebrated algorithm for contextual bandits with function approximation known as Optimistic Least Squares. Algorithm 1 presents the pseudo-code of the Optimistic Least Squares algorithm. In what follows,
@@ 5 Contextual Bandits with Unknown Variance
In the case where the variance is not known our contextual bandit algorithms work by estimating the cumulative variance up to constant multiplicative accuracy and use this estimator to build confidence sets as in Algorithms 1 and 2. In section 5.1 we describe how to successfully estimate the cumulative variance in contextual bandit problems and finally in section 5.3 we introduce Algorithm 3 that satisfy a regret guarantee whose dominating term scales with $\sqrt{\sum_{t=1}^{T}\sigma_{t}^{2}}$, and the low order term with $d_{\mathrm{eluder}}\cdot\log(|\mathcal{F}|)$.
@@ 5.2 Unknown-Variance Guarantees for Algorithm 2
Let $\delta\in(0,1)$. If $\delta/2$ is the input to Algorithm 2 and $\widehat{\sigma}_{t}^{2}$ estimators are computed by setting $\delta^{\prime}=\delta/2$, the regret of Algorithm 2 with modified confidence set sizes satisfies,
$\mathrm{Regret}(T)\leq\mathcal{O}\left(Bd_{\mathrm{eluder}}(\mathcal{F},B/T)\log^{2}(T)\log(T|\mathcal{F}|/\delta)+\sigma\sqrt{d_{\mathrm{eluder}}(\mathcal{F},B/T)\log(T|\mathcal{F}|/\delta)T}\right).$

### v3

@@ Abstract
Many works have developed no-regret algorithms for contextual bandits with function approximation, where the mean rewards over context-action pairs belong to a function class $\mathcal{F}$. Although there are many approaches to this problem, algorithms based on the principle of optimism, such as optimistic least squares have gained in importance. The regret of optimistic least squares scales as $\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathrm{eluder}}(\mathcal{F})\log(\mathcal{F})T}\right)$ where $d_{\mathrm{eluder}}(\mathcal{F})$ is a statistical measure of the complexity of the function class $\mathcal{F}$ known as eluder dimension. Unfortunately, even if the variance of the measurement noise of the rewards at time $t$ equals $\sigma_{t}^{2}$ and these are close to zero, the optimistic least squares algorithm’s regret scales with $\sqrt{T}$. In this work we are the first to develop algorithms that satisfy regret bounds for contextual bandits with function approximation of the form $\widetilde{\mathcal{O}}\left(\sigma\sqrt{\log(\mathcal{F})d_{\mathrm{eluder}}(\mathcal{F})T}+d_{\mathrm{eluder}}(\mathcal{F})\cdot\log(|\mathcal{F}|)\right)$ when the variances are unknown and satisfy $\sigma_{t}^{2}=\sigma$ for all $t$ and $\widetilde{\mathcal{O}}\left(d_{\mathrm{eluder}}(\mathcal{F})\sqrt{\log(\mathcal{F})\sum_{t=1}^{T}\sigma_{t}^{2}}+d_{\mathrm{eluder}}(\mathcal{F})\cdot\log(|\mathcal{F}|)\right)$ when the variances change at every time-step. These bounds generalize existing techniques for deriving second order bounds in contextual linear problems.
@@ 1 Introduction
The focus of many works, including this one is to flesh out the consequences of different modeling assumptions governing the relationship between the context, the action and the reward; for example by developing algorithms for scenarios where the reward is a linear function of an embedding of the context and action pair [6, 34, 11, 1]. This has lead to algorithms such as OFUL that can be used to derive bounds for contextual bandit problems with linear rewards [1]. Other works have considered scenarios that go beyond the linear case, where it is assumed the reward function over context action pair $x_{t},a_{t}$ is realized by an unknown function $f_{\star}$ belonging to a known function class $\mathcal{F}$ (which can be more complex than linear). Various adaptive learning procedures compatible with generic function approximation have been proposed for contextual bandit problems. Among these, we highlight two significant methods relevant to our discussion; the Optimistic Least Squares algorithm introduced by [35] and the SquareCB algorithm introduced by [14]. Both of these methods offer guarantees for cumulative regret. Specifically, the cumulative regret of Optimistic Least Squares scales with factors $\mathcal{O}(\sqrt{d_{\mathrm{eluder}}(\mathcal{F})\log(|\mathcal{F}|)})$, while the cumulative regret of Square CB scales as $\mathcal{O}\left(\sqrt{|\mathcal{A}|\log(|\mathcal{F}|)}\right)$, where $\mathcal{A}$ is the set of actions. The eluder dimension [^1] We formally introduce this quantity in Section 2. Here we use a simpler notation to avoid confusion. ($d_{\mathrm{eluder}}$) is a statistical complexity measure introduced by [35], that enables deriving guarantees for adaptive learning algorithms based on the principle of optimism in contextual bandits and reinforcement learning [26, 19, 32, 10].
@@
A recent work [18], published while our paper was under review, removes the assumptions made in [42, 41]. Here a summary of their results. When the variances $\sigma_{t}$ are revealed with the contexts, they show that for some function classes, any algorithm must incur a regret of $\Omega\left(\sqrt{\min(|\mathcal{A}|,d_{\mathrm{eluder}})\Lambda}+\min(d_{\mathrm{eluder}},\sqrt{|\mathcal{A}|T})\right)$, where $\Lambda=\sum_{t=1}^{T}\sigma^{2}$. They also propose an algorithm with an upper bound of $\mathcal{O}\left(\sqrt{|\mathcal{A}|\Lambda\log(|\mathcal{F}|)}+d_{\mathrm{eluder}}\log(|\mathcal{F}|)\right)$. In this setting, our techniques from Section 4 yield a refined bound of $\mathcal{O}\left(\sqrt{d_{\mathrm{eluder}}\Lambda\log(|\mathcal{F}|)}+d_{\mathrm{eluder}}\log(|\mathcal{F}|)\right)$ (see Theorem 4.5 for the special case where all variances are equal). For the setting where variances are not revealed with $x_{t}$ and may depend on the action $a_{t}$[18] derives the lower bound $\Omega\left(\min\left(\sqrt{d_{\mathrm{eluder}}\Lambda}+d_{\mathrm{eluder}},\sqrt{|\mathcal{A}|T}\right)\right)$. For the unknown fixed-variance case, our bounds match their lower bound (see Theorem 5.6). They also present an algorithm achieving an upper bound of $\mathcal{O}\left(d_{\mathrm{eluder}}\sqrt{\Lambda\log(|\mathcal{F}|)}+d_{\mathrm{eluder}}\log(|\mathcal{F}|)\right)$, which matches our result in Theorem 5.10.
@@ 3 Optimistic Least Squares
The algorithms we propose in this work are based on the optimism principle. This simple yet powerful algorithmic idea is the basis of a celebrated algorithm for contextual bandits with function approximation known as Optimistic Least Squares. Algorithm 1 presents the pseudo-code of the Optimistic Least Squares algorithm.
@@ 5 Contextual Bandits with Unknown Variance
In the case where the variance is not known our contextual bandit algorithms work by estimating the cumulative variance up to constant multiplicative accuracy and use this estimator to build confidence sets as in Algorithms 1 and 2. In section 5.1 we describe how to successfully estimate the cumulative variance in contextual bandit problems, in section 5.2 we show how to adapt a version of Algorithm 2 to the case of a single unknown variance and finally in section 5.3 we introduce Algorithm 3 that satisfies a regret guarantee whose dominating term scales with $d_{\mathrm{eluder}}\sqrt{\log(|\mathcal{F}|)\sum_{t=1}^{T}\sigma_{t}^{2}}$, and the low order term with $d_{\mathrm{eluder}}\cdot\log(|\mathcal{F}|)$.
@@ 5.2 Unknown-Variance Guarantees for Algorithm 2
Let $\delta\in(0,1)$. If $\delta/2$ is the input to Algorithm 2 and $\widehat{\sigma}_{t}^{2}$ estimators are computed by setting $\delta^{\prime}=\delta/2$. If $\sigma_{t}=\sigma$ for all $t\in\mathbb{N}$ the regret of Algorithm 2 with modified confidence set sizes satisfies,
$\mathrm{Regret}(T)\leq\mathcal{O}\left(\sigma\sqrt{d_{\mathrm{eluder}}(\mathcal{F},B/T)\log(T|\mathcal{F}|/\delta)T}+Bd_{\mathrm{eluder}}(\mathcal{F},B/T)\log^{2}(T)\log(T|\mathcal{F}|/\delta)\right).$